OpenStudy (anonymous):

in a rt.isosceles triangle,the lengths of the two non hypotenuse sides are designated as a.What is the perimeter of the triangle in terms of a?

OpenStudy (anonymous):

@amriju

OpenStudy (amriju):

use pythagoras to determine the hypotenuse....hyp=\[\sqrt{a ^{2}+a^{2}}\]...since both sides r of equal length a...so hyp=\[\sqrt{2}a\]...so we have the three sides..add them..dats ur perimeter=2a+\[\sqrt{2}a\]

OpenStudy (amriju):

did u get it..?

OpenStudy (anonymous):

yes thankyou

OpenStudy (anonymous):

the first term in a sequence is -5 and each subsequent term is 6 more than the immediately preceding term.What is the value of the 104th term?

OpenStudy (anonymous):

a104 = a + 103d = -5 + (103)*6

OpenStudy (anonymous):

in the sequence 1/8,1/4,1/2,..... each term after the first is equal to the previous terms a constant.What is the value of the 13th term?

OpenStudy (anonymous):

@amriju

OpenStudy (amriju):

u got to know geometric progression

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

solve it